Management is uncomfortable stating probabilities for the states of nature for a trade bill passing. Depending on its probability they will invest in one of three countries; A, B, or C.
Let p denote the probability of the bill passing.0/5 pts Question 10 Management is uncomfortable stating probabilities for the states of nature for a trade bill passing. Depe

Please explain to me how this answer was achieved. Thanks!
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

The expected value of :

A: 320*.65 + 160*(1-.65) = 264
B: 260*.65 + 210*(1-.65) = 242.5
C: 370*.65 + 0*(1-.65) = 240.5

A has the highest expected value, and therefore A should be selected